# EgoGaussian: Dynamic Scene Understanding from Egocentric Video with 3D Gaussian Splatting

### v1

@@ Abstract
Human activities are inherently complex, and even simple household tasks involve numerous object interactions. To better understand these activities and behaviors, it is crucial to model their dynamic interactions with the environment. The recent availability of affordable head-mounted cameras and egocentric data offers a more accessible and efficient means to understand dynamic human-object interactions in 3D environments. However, most existing methods for human activity modeling either focus on reconstructing 3D models of hand-object or human-scene interactions or on mapping 3D scenes, neglecting dynamic interactions with objects. The few existing solutions often require inputs from multiple sources, including multi-camera setups, depth-sensing cameras, or kinesthetic sensors. To this end, we introduce EgoGaussian, the first method capable of simultaneously reconstructing 3D scenes and dynamically tracking 3D object motion from RGB egocentric input alone. We leverage the uniquely discrete nature of Gaussian Splatting and segment dynamic interactions from the background. Our approach employs a clip-level online learning pipeline that leverages the dynamic nature of human activities, allowing us to reconstruct the temporal evolution of the scene in chronological order and track rigid object motion. Additionally, our method automatically segments object and background Gaussians, providing 3D representations for both static scenes and dynamic objects. EgoGaussian outperforms previous NeRF and Dynamic Gaussian methods in challenging in-the-wild videos and we also qualitatively demonstrate the high quality of the reconstructed models.  [^1] Project Website
@@ 1 Introduction
To reconstruct the dynamic scenes from an egocentric video, EgoGaussian first obtains hand-object segmentation using an off-the-shelf method and derives camera poses through structure-from-motion. By leveraging the natural trajectories of interactions, we partition the input video into static and dynamic clips. The static clips are used to reconstruct the background scenes and initialize the shapes of the object that will be interacted with. Subsequently, we refine the object’s shapes and track their motion through the dynamic clips. We empirically show that EgoGaussian achieves better reconstruction of dynamic scenes than previous NeRF and Dynamic Gaussian methods. We quantitatively evaluate our method on two in-the-wild egocentric video datasets following the evaluation protocol for novel-view synthesis. We also qualitatively demonstrate the high quality of the reconstructed scenes and the tracked object shapes and their motion.
Our main contributions can be summarized as follows: 1) We present a novel method that accurately reconstructs 3D scenes and dynamic object motion within them from RGB egocentric videos. 2) We leverage the dynamic nature of interactions that consist of transitions between static and dynamic phases, which facilitates the reconstruction of the static scenes, the object shapes, and the tracking of their motion. 3) Through both qualitative and quantitative evaluation, we demonstrate that our method outperforms previous approaches and provides better 4D reconstruction that captures the dynamic object interactions. Our model and code will be released upon acceptance.
@@ 2 Related Work
Hand-Object Segmentation. Many works have studied hand-object interaction in egocentric vision from different aspects. One significant area of focus is segmentation, specifically obtaining image segmentation masks of hands and the objects they hold. Ren et al. [49] proposed a motion-based approach to robustly segment both hand and object using optical flow and domain-specific cues from egocentric video. Concurrent with the emergence of deep neural networks-based hand-object segmentation is the scaling-up of egocentric data that includes pixel-level annotations and involves diverse daily activities [11, 12, 20]. VISOR[13] annotates videos from EPIC-KITCHENS[11, 12] dataset and provides masks for 67k hand-object relations covering 36 hours of videos. EgoHOS [74] further introduces the notion of a dense contact boundary to explicitly model the interaction and a context-aware compositional data augmentation technique to generate semantically consistent hand-object segmentation on out-of-distribution egocentric videos. Cheng et al. [9] produces a rich, unified 2D output of interaction by converting predicted bounding boxes to segments with Segment Anything (SAM) [30]. Our method takes egocentric videos with hand-object segmentation masks as input and creates dynamic 3D models.
@@
Hand-Object Reconstruction. Another highly related direction is to reconstruct the hand-object interaction, featuring 3D pose estimation for hands and objects. Recent works often jointly reconstruct hands and objects to favor physically plausible interactions [16, 33, 61, 69, 70, 78]. These approaches can be grouped into two categories. One assumes a known 3D object model and fits that model into 2D image [6, 10, 33, 55, 67]. For example, RHO [6] adapts an optimization-based approach that’s able to reconstruct hands and objects from single images in the wild, by leveraging 2D image cues and 3D contact priors to provide constraints. Recent works eliminate the need for a known 3D model and directly reconstruct 3d object shapes from the input [16, 48, 70]. However, they either require multiview input [48], specific hand-object interaction supervision [70], or can only reconstruct simple object shapes [16]. Current shape-agnostic methods struggle in in-the-wild scenarios [16, 78]. In contrast, our method does not require prior knowledge and obtains 3D object shapes through differentiable 3D Gaussian-based rendering.
@@
Static Scene Modeling. In the past few years, the domain of static scene modeling has garnered considerable attention. Mildenhall et al. [42] introduce the groundbreaking Neural Radiance Fields (NeRF), which utilizes a large Multilayer Perceptron (MLP) to represent 3D scenes and renders via volume rendering technique. However, their method queries the MLP at hundreds of points for each ray, resulting in slow training and rendering speed. Additionally, the original NeRF’s performance can diminish in scenes with highly dynamic elements due to its static, volumetric nature. Therefore, some subsequent works have tried to enhance the quality by (1) mitigating existing problems, such as aliasing [3, 4, 8] and reflection [24, 58] (2) incorporating image processing [41, 39] (3) employing per-image transient latent codes [40, 51], and (4) introducing supervision of expected depth with point clouds [14, 65]. There also exist some other follow-up works aiming to improve the speed, for example, by caching precomputed MLP results [27, 72], employing well-designed data structures [7, 54], removing the neural network [18], or utilizing multi-resolution hash encoding [43]. Yet, most of these methods still use ray marching, which involves sampling millions of points and slows down real-time rendering. Recently, Kerbl et al. [29] propose a different approach in the modeling and rendering of complex static 3D scenes - 3D Gaussian Splatting (3DGS). They model static scenes with Gaussians whose position, opacity, shape, and color are learned through a differentiable splatting-based renderer, achieving real-time rendering speed.
@@
Dynamic Scene Modeling. Motivated by the success of NeRF [42] in static scene modeling, numerous studies have adopted neural representations to model dynamic scenes. One strategy to extend 3D into 4D scenes is by using time stamps as an additional conditioning factor. [2, 64]. Another set of ’dynamic NeRF’ works [5, 17, 50, 52, 59] involves employing 4D space-time grid-based representations. Representing the 3D scene at a certain timestamp as a canonical space and then explicitly modeling deformation fields to warp 3D points into the canonical space is a common strategy as well [15, 34, 45, 47]. Another strategy is to combine the two approaches, using a conditional neural volume together with a deformation field [31, 46]. However, these methods all suffer from the same issues as static NeRFs, in that they require raymarching and despite advances in performance, still are not sufficiently fast for real-time rendering. Similarly, many dynamic extensions to 3D Gaussian splatting were also proposed [36, 37, 63, 68]. The most common approach is to learn for every timestep a set of deformations for each Gaussians. This can be done explicitly, or implicitly using a deformation which is evaluated for each Gaussian. This results in substantially faster training and rendering speed, with comparable levels of rendering quality. Although these methods result in decent quality renders, upon closer inspection all of them result in noticeably blurrier results than are possible with static reconstructions, especially when strong motion is involved.
@@ 3 Methodology
Figure 1 summarizes our method, EgoGaussian, for dynamic scene reconstruction from RGB egocentric videos. EgoGaussian first obtains camera poses and hand-object segmentation masks and the segmentation masks are further used to partition the videos into static and dynamic clips (Sec. 3.2). The static clips are used to initialize the static background and object shapes (Sec. 3.3), while the dynamic clips are used to track object motion and gradually refine their shapes (Sec. 3.4.
Figure 1: EgoGaussian Pipeline. Given an egocentric video input, our framework first estimates camera poses via structure-from-motion and obtains hand-object segmentation masks using an off-the-shelf approach. We also partition the video input into static and dynamic clips in the preprocessing step. The static clips are used to reconstruct the background scenes and initialize the shapes of the object that will be interacted with. Subsequently, we refine the object’s shapes and track their motion through the dynamic clips.
@@ 3.2 Data preprocessing
Camera pose estimation. For egocentric dataset where camera poses are already computed and made public, e.g., EPIC Fields [56] provides estimated camera poses for EPIC-KITCHENS [11], we will employ them directly; otherwise, we estimate camera poses ourselves. We will provide a frame filtering pipeline here to alleviate the computational load of running COLMAP on lengthy videos. COLMAP’s SfM also creates a sparse point cloud corresponding to the camera poses estimated, which we use as an initialization for 3D Gaussian Splatting.
Hand-object segmentation. We require pixel-level segmentation of hand-object interaction either provided [35] or estimated beforehand. The onset and offset frames of each hand-object interaction should also be estimated throughout the video. For the HOI4D dataset this is provided, whereas for the EPIC-KITCHENS dataset we use EgoHOS [74] to generate hand masks and Track-Anything model [66] for object masks and human body masks. Furthermore, these masks are dilated by 2 pixels for better robustness.
@@
Video partitioning. We then partition the egocentric video into static and dynamic clips according to the onset and offset of interactions. We define static clips as ones where only the actor’s hands or body are moving, but objects are all static, while dynamic clips contain both actor and object motion.
3.3 Static clip reconstruction
As some other works have pointed out, 3D Gaussian Splatting tends to overfit to training views by generating excessive floaters when there are scene inconsistencies among 3D views [22]. In order to eliminate such inconsistencies, any objects that move at all in the scene should be identified and masked out to provide a purely static scene reconstruction. Currently moving objects can be identified using the masks mentioned in the previous section. We explain in the following sections how we also identify objects that have moved or will move, and we provide an illustration of this process in Figure 2
Figure 2: Static reconstruction pipeline. Given a static clip partitioned from egocentric video, we adopt a standard 3D Gaussian Splatting training schema first. Based on the trained Gaussians, we then train an object identity variable with respect to 5 object masks, so that object Gaussians can be detached from the background Gaussians.
@@ 3.3 Static clip reconstruction
Initial Static training. We first train a static version of the scene with each static clip. To do so, we use a set of $M$ observations/frames from this clip $S_{i}=\{\mathbf{I}_{s_{j}},\mathbf{H}_{s_{j}},\theta_{s_{j}}|$  $j=1,\dots,M\}$, where $\mathbf{I}_{j}$ is an input RGB egocentric frame, $\mathbf{H}_{j}$ is the binary hand/body mask where pixel value $=0$ represents body part and pixel value $=1$ is for rest of the frame, and $\theta_{j}$ is the corresponding camera parameters for frame $j$. We then follow the similar optimization pipeline as the original 3DGS, including pruning and densification. We use a masked version of their loss function:
@@
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}\left(\mathbf{I}_{\text{input}},\mathbf{I}_{\text{render}}\right)+\lambda\mathcal{L}_{\text{D-SSIM }}\left(\mathbf{I}_{\text{input}},\mathbf{I}_{\text{render}}\right),$ (1)
@@
with the gradients zeroed out according to the hand/body mask $\mathbf{H}$. Similar to SuGaR [23], after around 30K iterations, we append an additional entropy loss on the opacity $\alpha$ of Gaussians, i.e.
$\mathcal{L}_{\text{entropy}_{\alpha}}=-\alpha\log(\alpha)-(1-\alpha)\log(1-\alpha),$ (2)
@@
as a way to enforce Gaussians to be either fully transparent or completely opaque and train for another 10K iterations while disabling pruning and densification. Instead, we prune the transparent Gaussians once at the end of this phase of training.
Object identity training. This produces a set of 3D Gaussians $\mathcal{G}$ reconstructing the scene captured by this static clip, which includes both the static background and any objects that might move during dynamic portions of the video. As previously mentioned, we require masks not only for currently moving objects, but objects that move at any time in the video. As these objects are static and not segmented during static clips, we instead use the masks from the closest adjacent dynamic frame, and copy them forwards and backwards in time. For example, for a static clip $S_{1}$ with $M$ frames between two dynamic ones where different objects are moved, masks for two different objects A and B should be determined for the first and last 5 frames of this clip respectively: $\{\tilde{\mathbf{O}}_{A,1},\dots\tilde{\mathbf{O}}_{A,5}\}$, $\{\tilde{\mathbf{O}}_{B,M-4},\dots\tilde{\mathbf{O}}_{B,M}\}$. Although this assumes that the camera moves minimally during these transitions, this works well in practice for all tested settings. An additional trainable parameter of label $l$ is then attached to each Gaussian and initialized to a very small value. This label can then be rendered similar to the RGB value as: $L(\mathbf{x}_{p})=\sum_{i\in\mathcal{N}}l_{i}\alpha_{i}\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right)$. This produces a segmentation mask $L_{render}$, upon which we can apply a binary cross entropy loss using the dynamic object masks $\tilde{\mathbf{O}}$ we previous discussed:
$\mathcal{L}_{\text{BCEWithLogitsLoss}}=-\left[(1-\tilde{\mathbf{O}})\cdot\log(\sigma(\mathbf{I}_{\text{render, obj}}))+\tilde{\mathbf{O}}\cdot\log(1-\sigma(\mathbf{I}_{\text{render, obj}})))\right],$ (3)
where $\sigma(X)=\frac{1}{1+e^{-X}}$ is the sigmoid function. This allows us to separate the Gaussians into object and background Gaussians based on thresholding $\sigma(l)$ and can be rendered to compute movable object masks.
Combination of static and dynamic clips. The previously mentioned static scene still includes objects that are only static for some parts of the video. However, parts of the background scene are obscured by these objects, and only become visible during dynamic clips. We therefore retrain the static scene using the movable object masks in order to mask out objects that move during any time in the video. This allows us to train a more complete static scene which includes portions of the background that are only visible during dynamic parts of the video.
3.4 Dynamic object modeling
Object pose estimation. The pre-trained and segmented set of Gaussians $\mathcal{G}_{\text{obj}}$ can be used not only to mask out the objects for purpose of background reconstruction, but also as an initial estimate of the object appearance. Then, drawing inspiration from the previous hand-object reconstruction works that use a differentiable renderer such as [6, 73], we estimate the object pose for each video frame in the dynamic clip. Specifically, we estimate for each frame $f_{j}$ a corresponding pose $\mathbf{p}_{j}$. We further decompose the pose $\mathbf{p}_{j}$ into a 3D translation vector $\mathbf{t}_{j}$ and a rotation matrix $\mathbf{R}_{j}$. Unlike previous dynamic Gaussian Splatting methods [63, 37], our method applies one set of transformation parameters to the whole collection as a whole, treating it as a single rigid object.
To ensure the transformation is rigid, any estimated $3\times 3$ rotation matrix must be $\mathbf{R}_{i}\in SO(3)$. We optimize the rotation $\mathbf{R}_{j}$ using the 6D continuous rotation representation proposed by [77] and translation $\mathbf{t}_{j}$. Hence for each input frame $f_{j}$, assuming we have the Gaussians representing the object’s state in the previous frame, $\mathcal{G}_{\text{obj},j-1}$, we apply the current trainable translation and rotation parameters to it: $\mathbf{X}_{\mathcal{G}_{\text{obj},j}}=\mathbf{X}_{\mathcal{G}_{\text{obj},j-1}}\cdot g(\tilde{\mathbf{R}}_{j})+\mathbf{t}_{j}$, where $\mathbf{X}_{\mathcal{G}_{\text{obj},j}}$ is the 3D coordinates of Gaussians $\mathcal{G}_{\text{obj},j}$ corresponding to the object’s state at frame $j$, $g(\cdot)$ is a function defined in [77] that transforms the 6D representation of rotation to a standard $3\times 3$ rotation matrix.
In order to prioritize learning the object pose, during this stage the learning rate on the parameters of the Gaussians such as position and color are all lowered by a factor of 10. In order to better constrain the shape of the object and prevent it from extending to outside of the mask, we also apply a silhouette loss onto the computed alpha value, which can be computed using the following equation: $A(\mathbf{x}_{p})=\sum_{i\in\mathcal{N}}\alpha_{i}\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right)$ which effectively equates the RGB rendering equation without color. Our final object loss is then:
$\mathcal{L}_{\text{obj}}=\mathcal{L}_{1}\left(\mathbf{I}_{\text{obj}},\mathbf{I}_{\text{render},\mathcal{G}_{\text{obj}}}\right)+\lambda\mathcal{L}_{2}\left(1-\mathbf{O},\mathbf{A}_{\text{render},\mathcal{G}_{\text{obj}}}\right),$ (4)
where $\mathbf{I}_{\text{obj}}$ is cropped from $\mathbf{I}_{\text{input}}$ with the object mask $\mathbf{O}$, $\mathbf{I}_{\text{render},\mathcal{G}_{\text{obj}}}$ is rendered from $\mathcal{G}_{\text{obj}}$ so it contains the object only and black background, $\mathbf{A}_{\text{render},\mathcal{G}_{\text{obj}}}$ is the rendered alpha. We experimentally observe that $0.5$ is a suitable $\lambda$ value.
As we optimize the time-dependent pose parameters $\mathbf{t}_{j}$ and $\tilde{\mathbf{R}}_{j}$ one frame at a time, the Gaussians can easily overfit to the current frame. To address this, we train not only on the current frame, instead for every training iteration we train either on the current frame or a random previous frame with a probability of $0.5$. A problem is that during the Gaussian splatting optimization process, the opacity is frequently set to zero in order to prune floaters. However, this would produce a very noisy signal for the pose optimization. As such, we instead alternate between optimizing the rigid object pose, and densifying/pruning the Gaussians. In practice, for every dynamic frame, we first train for 4k iterations only optimizing the object pose without pruning or densification. We then freeze the object pose, and train another 4k iterations to better incorporate any new information from the object pose, and finally train another 4k iterations again only the pose without densification or pruning. For all 12k iterations, all Gaussian parameters such as color and position are continuously optimized. After iterating through the whole dynamic clip with $M$ frames, we obtain a coarse object pose for each frame $P=\{\mathbf{t}_{j},\mathbf{R}_{j}|j=1,\ldots,M\}$. Finally, we perform one final round of training on all frames jointly, with 6k iterations of pose estimation, 6k iterations of pruning/densifications, and finally another 6k iterations of pose estimation. This ensures that our object model is more equally fit onto all frames, rather than focused on the last seen ones. We provide an illustration of this process in Figure 3

### v2

@@ Abstract
Human activities are inherently complex, often involving numerous object interactions. To better understand these activities, it is crucial to model their interactions with the environment captured through dynamic changes. The recent availability of affordable head-mounted cameras and egocentric data offers a more accessible and efficient means to understand human-object interactions in 3D environments. However, most existing methods for human activity modeling neglect the dynamic interactions with objects, resulting in only static representations. The few existing solutions often require inputs from multiple sources, including multi-camera setups, depth-sensing cameras, or kinesthetic sensors. To this end, we introduce EgoGaussian, the first method capable of simultaneously reconstructing 3D scenes and dynamically tracking 3D object motion from RGB egocentric input alone. We leverage the uniquely discrete nature of Gaussian Splatting and segment dynamic interactions from the background, with both having explicit representations. Our approach employs a clip-level online learning pipeline that leverages the dynamic nature of human activities, allowing us to reconstruct the temporal evolution of the scene in chronological order and track rigid object motion. EgoGaussian shows significant improvements in terms of both dynamic object and background reconstruction quality compared to the state-of-the-art. We also qualitatively demonstrate the high quality of the reconstructed models. Video demonstrations, code, and datasets will be available at https://zdwww.github.io/egogs.github.io.
Figure 1: EgoGaussian Pipeline. Given an egocentric video input, our framework first partitions the video input into static and dynamic clips and obtains hand-object segmentation masks using off-the-shelf approaches. The static clips are then used to reconstruct the background scenes and initialize the shapes of the object that will be interacted with using 3D Gaussians. Subsequently, we refine the object’s shapes and track their motion through the dynamic clips. EgoGaussian enables a high-quality 4D reconstruction of dynamic scenes with explicit representations that accurately capture the object interactions.
@@ 1 Introduction
To reconstruct the dynamic scenes from an egocentric video, EgoGaussian first obtains hand-object segmentation using an off-the-shelf method and derives camera poses through structure-from-motion. By leveraging the natural trajectories of interactions, we partition the input video into static and dynamic clips. The static clips are used to reconstruct the background scenes and initialize the shapes of the object that will be interacted with. Subsequently, we refine the object’s shapes and track their motion through the dynamic clips. We empirically show that EgoGaussian achieves better reconstruction of dynamic scenes than the state-of-the-art. We quantitatively evaluate our method on two in-the-wild egocentric video datasets following the evaluation protocol for novel-view synthesis. We also qualitatively demonstrate the high quality of the reconstructed scenes and the tracked object shapes and their motion.
Our main contributions can be summarized as follows:
  • We present a novel method that accurately reconstructs 3D scenes and dynamic object motion within them from RGB egocentric videos.
  • We leverage the dynamic nature of interactions that consist of transitions between static and dynamic phases, which facilitates the reconstruction of the static scenes, the object shapes, and the tracking of their motion.
  • Through both qualitative and quantitative evaluation, we demonstrate that our method outperforms previous approaches and provides better 4D reconstruction that captures the dynamic object interactions.
@@ 2 Related Work
Hand-Object Segmentation. Many works have studied hand-object interaction in egocentric vision from different aspects. One significant area of focus is segmentation, specifically obtaining image segmentation masks of hands and the objects they hold. Ren et al. [48] proposed a motion-based approach to robustly segment both hand and object using optical flow and domain-specific cues from egocentric video.
@@
Concurrent with the emergence of deep neural networks-based hand-object segmentation is the scaling-up of egocentric data that includes pixel-level annotations and involves diverse daily activities [11, 12, 19]. VISOR[13] annotates videos from EPIC-KITCHENS[11, 12] dataset and provides masks for 67k hand-object relations covering 36 hours of videos. EgoHOS [74] further introduces the notion of a dense contact boundary to explicitly model the interaction and a context-aware compositional data augmentation technique to generate semantically consistent hand-object segmentation on out-of-distribution egocentric videos. Cheng et al. [9] produces a rich, unified 2D output of interaction by converting predicted bounding boxes to segments with Segment Anything (SAM) [29]. Our method takes egocentric videos with hand-object segmentation masks as input and creates dynamic 3D models.
@@
Hand-Object Reconstruction. Another highly related direction is to reconstruct the hand-object interaction, featuring 3D pose estimation for hands and objects. Recent works often jointly reconstruct hands and objects to favor physically plausible interactions [16, 32, 61, 70, 71, 78]. These approaches can be grouped into two categories. One assumes a known 3D object model and fits that model into 2D image [6, 10, 32, 56, 67]. For example, RHO [6] adapts an optimization-based approach that’s able to reconstruct hands and objects from single images in the wild, by leveraging 2D image cues and 3D contact priors to provide constraints.
Recent works eliminate the need for a known 3D model and directly reconstruct 3d object shapes from the input [16, 47, 71]. However, they either require multiview input [47], specific hand-object interaction supervision [71], or can only reconstruct simple object shapes [16]. Current shape-agnostic methods struggle in in-the-wild scenarios [16, 78]. In contrast, our method does not require prior knowledge and obtains 3D object shapes through differentiable 3D Gaussian-based rendering.
@@
Static Scene Modeling. In the past few years, the domain of static scene modeling has garnered considerable attention. Mildenhall et al. [40] introduce the groundbreaking Neural Radiance Fields (NeRF), which utilizes a large Multilayer Perceptron (MLP) to represent 3D scenes and renders via volume rendering technique. However, their method queries the MLP at hundreds of points for each ray, resulting in slow training and rendering speed. Additionally, the original NeRF’s performance can diminish in scenes with highly dynamic elements due to its static, volumetric nature. Therefore, some subsequent works have tried to enhance the quality by (1) mitigating existing problems, such as aliasing [3, 4, 8] and reflection [23, 58] (2) incorporating image processing [41, 38] (3) employing per-image transient latent codes [39, 52], and (4) introducing supervision of expected depth with point clouds [14, 65]. There also exist some other follow-up works aiming to improve the speed, for example, by caching precomputed MLP results [26, 73], employing well-designed data structures [7, 55], removing the neural network [18], or utilizing multi-resolution hash encoding [42]. Yet, most of these methods still use ray marching, which involves sampling millions of points and slows down real-time rendering.
Recently, Kerbl et al. [28] propose a different approach in the modeling and rendering of complex static 3D scenes - 3D Gaussian Splatting (3DGS). They model static scenes with Gaussians whose position, opacity, shape, and color are learned through a differentiable splatting-based renderer, achieving real-time rendering speed.
@@
Dynamic Scene Modeling. Motivated by the success of NeRF [40] in static scene modeling, numerous studies have adopted neural representations to model dynamic scenes. One strategy to extend 3D into 4D scenes is by using time stamps as an additional conditioning factor. [2, 64]. Another set of ’dynamic NeRF’ works [5, 17, 49, 53, 59] involves employing 4D space-time grid-based representations. Representing the 3D scene at a certain timestamp as a canonical space and then explicitly modeling deformation fields to warp 3D points into the canonical space is a common strategy as well [15, 34, 44, 46]. Another strategy is to combine the two approaches, using a conditional neural volume together with a deformation field [30, 45]. However, these methods all suffer from the same issues as static NeRFs, in that they require raymarching and despite advances in performance, still are not sufficiently fast for real-time rendering.
Similarly, many dynamic extensions to 3D Gaussian splatting were also proposed [35, 36, 63, 68]. The most common approach is to learn for every timestep a set of deformations for each Gaussians. This can be done explicitly, or implicitly using a deformation which is evaluated for each Gaussian. This results in substantially faster training and rendering speed, with comparable levels of rendering quality. Although these methods result in decent quality renders, upon closer inspection all of them result in noticeably blurrier results than are possible with static reconstructions, especially when strong motion is involved.
@@ 3 Methodology
Figure 1 summarizes our method, EgoGaussian, for dynamic scene reconstruction from RGB egocentric videos. Our central idea is to identify dynamic objects and train them separately from the background. To do so, we develop a framework that fully integrates the dynamic characteristics of object interactions by temporally segmenting the video into static and dynamic clips and applying existing hand-object interaction modeling techniques to achieve 2D spatial distinction. Specifically, EgoGaussian first obtains camera poses and hand-object segmentation masks and the segmentation masks are further used to partition the videos into static and dynamic clips (Sec. 3.2). The static clips are used to initialize the static background and object shapes (Sec. 3.3), while the dynamic clips are used to track object motion and gradually refine their shapes (Sec. 3.4).
@@ 3.2 Data Preprocessing
As pointed out by previous work [21], 3D-GS tends to overfit to training views and generate excessive floaters when there are scene inconsistencies among 3D views. In order to eliminate such inconsistencies, our idea is to identify any objects that move at all in the scene and separate them from the static background. In this preprocessing step, we adopt an initialisation method similar to other 3D-GS approaches and process the input to separate dynamic contents from static ones.
@@
Initialisation for 3D-GS. Following existing 3D-GS methods, we first use COLMAP [50, 51] to estimate camera poses. COLMAP’s SfM also creates a sparse point cloud corresponding to the camera poses estimated, and we use them as an initialization for 3D Gaussian Splatting. For egocentric dataset where camera poses are available, e.g., EPIC Fields [57] provides estimated camera poses for EPIC-KITCHENS [11], we employ them directly.
@@
Separation of Dynamic and Static. To separate dynamic and static information in 2D frames, we use off-the-shelf approaches to obtain segmentation masks of hand-object interaction. Specifically, we use EgoHOS [74] to generate hand masks and Track-Anything model [66] for object masks and human body masks. Furthermore, these masks are dilated by 2 pixels for better robustness. The onset and offset frames of each hand-object interaction are estimated throughout the video. We then partition the egocentric video along the temporal axis into static and dynamic clips according to the onset and offset of interactions. We define static clips as ones where only the actor’s hands or body are moving, but objects are all static, while dynamic clips contain both actor and object motion.
3.3 Static Reconstruction
Given a static video clip, our goal is to obtain a 3D reconstruction that distinguishes between the static background and dynamic objects involved in interactions within the dynamic clips. We begin by training a static representation of the scene and then identify dynamic objects using information extracted from the adjacent dynamic clips.
@@ 3.3 Static Reconstruction
Initial Static Reconstruction. We have a set of $T$ observations/frames from the static clip $S=\{\mathbf{I}_{t},\mathbf{M}_{\text{body},t},\theta_{t}|$  $t=1,\dots,T\}$ as input, where $\mathbf{I}_{t}$ is an input RGB egocentric frame, $\mathbf{M}_{\text{body},t}$ is the binary hand/body segmentation mask where pixel value $=0$ represents body part and pixel value $=1$ is for rest of the frame, and $\theta_{t}$ is the corresponding camera parameters for frame $t$. We follow a similar optimization pipeline as the original 3DGS [28], including both pruning and densification but use a masked version of the loss function:
@@
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}\left(\mathbf{I}_{\text{input}},\mathbf{I}_{\text{render}}\right)+\lambda\mathcal{L}_{\text{D-SSIM }}\left(\mathbf{I}_{\text{input}},\mathbf{I}_{\text{render}}\right),$
@@
with the gradients zeroed out according to the mask $\mathbf{M}_{\text{body}}$. Similar to SuGaR [22], after around 30K iterations, we append an additional entropy loss on the opacity $\alpha$ of Gaussians, i.e.
$\mathcal{L}_{\text{entropy}_{\alpha}}=-\alpha\log(\alpha)-(1-\alpha)\log(1-\alpha),$
@@
as a way to enforce Gaussians to be either fully transparent or completely opaque and train for another 10K iterations while disabling pruning and densification. Instead, we prune the transparent Gaussians once at the end of this phase of training. This produces a set of 3D Gaussians $\mathcal{G}$ reconstructing the scene captured by this static clip, which includes both the static background and any objects that might move during dynamic portions of the video.